OpenStudy (happykiddo):

Which of the following is a solution to the equation x^2 – 6x + 5 = 0? A. x = -5 B. x = -1 C. x = 1/5 D. x = 5

OpenStudy (happykiddo):

i know its D but how to get i dont know

OpenStudy (anonymous):

break it up into (x-1)(x-5) = 0

OpenStudy (anonymous):

solve for x-1 = 0 and x-5=0

OpenStudy (happykiddo):

i know i factored but it gives me two solutions 5 and 1

OpenStudy (happykiddo):

there is only supposed to be one (which is 5), as shown on the answer key

OpenStudy (precal):

2nd power means you are looking for 2 solutions. Which implies one is given and not all......

OpenStudy (happykiddo):

could you give me an example of what you mean precal

OpenStudy (precal):

this quadratic has 2 solutions because the function crossed the x axis in 2 places. your multiple choice only listed one of the 2 solutions